OpenStudy (willie579):

Tyrone's bank account had a balance of -$73.59 last week. His paycheck of $269.77 was deposited this week. If there have been no other transactions, which of the following is true? A.)Tyrone's current account balance is $196.18. B.)Tyrone's current account balance is -$196.18. C.)Tyrone's current account balance is $343.36. D.)Tyrone's current account balance is -$343.36.

OpenStudy (koikkara):

hai @Willie579 think like this... you had to give me 10 bucks. Later you got a lottery of 50 bucks. then how much amount is remaining with you ?

OpenStudy (willie579):

40?

OpenStudy (koikkara):

Excellent... how you found it ?

OpenStudy (willie579):

Just took 50 - 10. =/

OpenStudy (koikkara):

$269.77 -$73.59 = \(?\)

OpenStudy (willie579):

A.

OpenStudy (koikkara):

You did quiet well, Congrats

OpenStudy (willie579):

I have some more questions. :/

OpenStudy (willie579):

http://prntscr.com/98o5ag

OpenStudy (koikkara):

cool..

OpenStudy (koikkara):

here, the measurement hints looks like 1/5 of a foot long= (1 / 5) * (1 foot) = 0.2 feet (213⁄32 inches) 3/5 of a foot long =(3 / 5) * (1 foot) = 0.6 feet (713⁄64 inches)

OpenStudy (willie579):

I got it. :)

OpenStudy (willie579):

http://prntscr.com/98o7xo

OpenStudy (willie579):

* http://prntscr.com/98o8yw

OpenStudy (koikkara):

Hint: 1 Batch= 2 t-spoon 4 Batch= ?

OpenStudy (willie579):

8.

OpenStudy (willie579):

So the answer is A.

OpenStudy (koikkara):

no.... your screen shot is wrong... it should be Hint: 1 Batch= \(1\2\) t-spoon 4 Batch= ?

OpenStudy (willie579):

yeah, the screenshot was wrong so i posted a new one. my bad. :/

OpenStudy (koikkara):

please check \(?\2\)

OpenStudy (willie579):

But I know that the answer is A. Because 1/2 x 4 = 2.

OpenStudy (willie579):

And 1/2 = 1.5. 1.5 x 4 = 2/

OpenStudy (koikkara):